# VCR: Visual Caption Restoration

### v3

@@ Appendix A Additional evaluation results on first 100 and 500 test cases
Table 4: Results of various open-source and closed-source vision language models on the VCR task using the first 100 test cases. FT = finetuned on 16,000 samples from the VCR-wiki training set. The best results among the finetuned models are underlined while the best results among the models without finetuning are highlighted in bold. Subscripts are standard deviations obtained from Bootstrap.
@@
CogVLM2-FT 19B $\underline{94.08_{0.2}}$ $\underline{94.67_{0.26}}$0.59 $\underline{98.03_{0.07}}$ $\underline{98.22_{0.03}}$0.2
DeepSeek-VL 1.3B $19.53_{0.69}$ $26.04_{1.47}$6.51 $43.73_{0.18}$ $48.03_{0.16}$4.3
@@
Llama-3.2-Vision 11B $79.88_{2.96}$ $68.64_{3.75}$ 11.24 $90.88_{1.56}$ $82.91_{2.11}$ 7.97
Llama-3.2-Vision 90B $79.29_{3.14}$ $71.01_{3.36}$ 8.28 $87.81_{2.00}$ $83.17_{2.30}$ 4.64
@@
Qwen2-VL 7B $\bm{90.53_{2.25}}$ $\bm{96.45_{1.39}}$5.92 $\bm{94.28_{1.54}}$ $\bm{98.82_{0.49}}$4.54
@@
CogVLM2-FT 19B $\underline{75.74_{0.72}}$ $\underline{67.46_{0.64}}$ 8.28 $\underline{90.6_{0.13}}$ $\underline{84.26_{0.08}}$ 6.34
DeepSeek-VL 1.3B $0.0_{0.0}$ $0.0_{0.0}$ 0 $11.17_{0.03}$ $10.88_{0.06}$ 0.29
@@
Llama-3.2-Vision 11B $10.65_{2.33}$ $7.69_{2.04}$ 2.96 $33.50_{2.28}$ $26.80_{2.03}$ 6.70
Llama-3.2-Vision 90B $13.02_{2.59}$ $15.38_{2.75}$2.37 $36.80_{2.36}$ $39.85_{2.52}$3.06
@@
Qwen2-VL 7B $\bm{75.74_{3.32}}$ $\bm{73.96_{3.55}}$ 1.78 $\bm{85.91_{2.17}}$ $\bm{85.83_{2.04}}$ 0.08
@@
CogVLM2-Chinese-FT 19B $\underline{66.49_{0.74}}$ $\underline{67.55_{0.73}}$1.06 $\underline{79.48_{0.17}}$ $\underline{81.78_{0.09}}$2.3
DeepSeek-VL 1.3B $0.0_{0.0}$ $0.0_{0.0}$ 0 $6.69_{0.07}$ $2.92_{0.02}$ 3.78
@@
Qwen2-VL 7B $\bm{67.55_{3.34}}$ $\bm{73.40_{3.21}}$5.85 $\bm{84.63_{1.80}}$ $\bm{87.12_{1.76}}$2.49
@@
CogVLM2-Chinese-FT 19B $\underline{46.81_{0.32}}$ $\underline{46.28_{0.49}}$ 0.53 $\underline{66.85_{0.39}}$ $\underline{69.79_{0.12}}$2.95
DeepSeek-VL 1.3B $0.0_{0.0}$ $0.0_{0.0}$ 0 $6.5_{0.03}$ $4.16_{0.03}$ 2.34
@@
Yi-VL 6B $0.0_{0.0}$ $0.0_{0.0}$ 0 $4.15_{0.06}$ $2.38_{0.04}$ 1.77
Table 5: Results of various open-source and closed-source vision language models on the VCR task using the first 500 test cases. FT = finetuned on 16,000 samples from the VCR-wiki training set. The best results among the finetuned models are underlined while the best results among the models without finetuning are highlighted in bold. Subscripts are standard deviations obtained from Bootstrap.
@@
CogVLM2-FT 19B $\underline{92.8_{0.06}}$ $\underline{92.67_{0.13}}$ 0.12 $\underline{97.51_{0.24}}$ $\underline{97.45_{0.07}}$ 0.06
DeepSeek-VL 1.3B $21.86_{0.17}$ $30.68_{0.3}$8.82 $45.4_{0.33}$ $52.02_{0.73}$6.62
@@
Llama-3.2-Vision 11B $79.25_{1.40}$ $66.46_{1.63}$ 12.80 $89.98_{0.77}$ $80.91_{1.06}$ 9.08
Llama-3.2-Vision 90B $80.87_{1.37}$ $71.55_{1.54}$ 9.32 $89.63_{0.85}$ $84.34_{0.98}$ 5.29
@@
Qwen2-VL 7B $\bm{90.06_{1.07}}$ $\bm{94.53_{0.82}}$4.47 $\bm{93.77_{0.76}}$ $\bm{97.80_{0.34}}$4.03
@@
CogVLM2-FT 19B $\underline{75.9_{0.13}}$ $\underline{65.22_{0.18}}$ 10.68 $\underline{89.75_{0.14}}$ $\underline{82.71_{0.27}}$ 7.04
DeepSeek-VL 1.3B $0.37_{0.02}$ $0.12_{0.01}$ 0.25 $11.42_{0.09}$ $11.41_{0.22}$ 0.01
@@
Llama-3.2-Vision 11B $13.91_{1.25}$ $7.33_{0.94}$ 6.58 $35.78_{1.14}$ $26.14_{0.94}$ 9.64
Llama-3.2-Vision 90B $15.16_{1.28}$ $12.17_{1.13}$ 2.98 $37.57_{1.13}$ $35.14_{1.04}$ 2.43
@@
Qwen2-VL 7B $\bm{76.27_{1.49}}$ $\bm{75.65_{1.45}}$ 0.62 $\bm{86.56_{1.00}}$ $\bm{86.77_{0.93}}$0.22
@@
CogVLM2-Chinese-FT 19B $\underline{63.97_{0.55}}$ $\underline{62.67_{0.17}}$ 1.3 $\underline{79.71_{0.41}}$ $\underline{79.22_{0.47}}$ 0.49
DeepSeek-VL 1.3B $0.0_{0.0}$ $0.0_{0.0}$ 0 $6.1_{0.1}$ $3.25_{0.05}$ 2.85
@@
Qwen2-VL 7B $\bm{61.08_{1.52}}$ $\bm{68.16_{1.46}}$7.09 $\bm{78.38_{0.96}}$ $\bm{83.48_{0.84}}$5.10
@@
CogVLM2-Chinese-FT 19B $\underline{42.51_{0.32}}$ $\underline{45.91_{0.23}}$3.39 $\underline{65.79_{0.24}}$ $\underline{69.46_{0.46}}$3.68
DeepSeek-VL 1.3B $0.0_{0.0}$ $0.0_{0.0}$ 0 $6.87_{0.09}$ $3.53_{0.07}$ 3.33
@@
Yi-VL 6B $0.0_{0.0}$ $0.0_{0.0}$ 0 $3.95_{0.05}$ $2.08_{0.09}$ 1.87
Appendix B More relationship between VCR and other benchmarks
@@
Appendix C Information of models evaluated
We provide the specifications of the evaluated models and a link to each model’s weights in Table 6.
Table 6: Model specifications
Model name Model size Open-sourced
Claude 3 Opus -
[…]
2613
Claude 3.5 Sonnet -
[…]
2613
Gemini 1.5 Pro -
[…]
2613
GPT-4 Turbo -
[…]
2613
GPT-4o -
[…]
2613
GPT-4V -
[…]
2613
Qwen-VL-Max -
[…]
2613
Reka Core -
[…]
2613
Cambrian-1  [^5] https://huggingface.co/nyu-visionx/cambrian-34b 34B $\checkmark$
CogVLM2  [^6] https://huggingface.co/THUDM/CogVLM2-Llama3-chat-19B 19B $\checkmark$
CogVLM2-Chinese  [^7] https://huggingface.co/THUDM/cogvlm2-llama3-Chinese-chat-19B 19B $\checkmark$
DeepSeek-VL  [^8] https://huggingface.co/deepseek-ai/deepseek-vl-1.3b-chat 1.3B $\checkmark$
DeepSeek-VL  [^9] https://huggingface.co/deepseek-ai/deepseek-vl-7b-chat 7B $\checkmark$
Idefics3  [^11] https://huggingface.co/HuggingFaceM4/Idefics3-8B 8B $\checkmark$
InternLM-XComposer2-VL  [^12] https://huggingface.co/internlm/internlm-xcomposer2-vl-7b 7B $\checkmark$
InternLM-XComposer2-VL-4KHD  [^13] https://huggingface.co/internlm/internlm-xcomposer2-4khd-7b 7B $\checkmark$
InternVL-V2  [^14] https://huggingface.co/OpenGVLab/InternVL2-40B 40B $\checkmark$
Llama-3.2-Vision  [^15] https://huggingface.co/meta-llama/Llama-3.2-11B-Vision-Instruct 11B $\checkmark$
Llama-3.2-Vision  [^16] https://huggingface.co/meta-llama/Llama-3.2-90B-Vision-Instruct 90B $\checkmark$
MiniCPM-V2.5  [^17] https://huggingface.co/OpenBMB/MiniCPM-Llama3-V-2_5 8B $\checkmark$
Qwen2-VL  [^21] https://huggingface.co/Qwen/Qwen2-VL-7B-Instruct 7B $\checkmark$
Yi-VL [^22] https://huggingface.co/01-ai/Yi-VL-34B 34B $\checkmark$
Yi-VL [^23] https://huggingface.co/01-ai/Yi-VL-6B 6B $\checkmark$
Appendix D Potential QA
@@
    Include VCR in VLM pretraining: Just as OCR parsing tasks are often included in pretraining to improve OCR performance, researchers could consider incorporating VCR tasks during pretraining. We have released vcr_transform.py [^24] https://github.com/tianyu-z/VCR/blob/main/src/build_dataset/vcr_transform.py to facilitate this process, making it as straightforward as data augmentation.
@@
    Chain-of-Thought (CoT) Methods: Researchers could explore multi-modality pipelines based on CoT techniques to improve existing VLMs on VCR tasks Chen et al. (2024a); Zhang et al. (2023). Although a model might not initially focus on the correct visual area (e.g., pixel-level hints in the TEI), CoT-based techniques could help refine its focus over successive rounds.
@@
Table 7: Performance Comparison of Base Models

### v4

@@ Appendix D Additional evaluation results on first 100 and 500 test cases
Table 8: Results of various open-source and closed-source vision language models on the VCR task using the first 100 test cases. FT = fine-tuned on 16,000 samples from the VCR-wiki training set. We label the best result of each setting and metric with bold fonts, and the best open-souce model with underline. Subscripts are standard deviations obtained from Bootstrap. For more latest models’ results, please visit our GitHub repository.
@@
CogVLM2-FT 19B $94.08_{0.2}$ $94.67_{0.26}$0.59 $98.03_{0.07}$ $98.22_{0.03}$0.2
@@
InternVL-V2 76B $88.17_{2.48}$ $92.31_{2.05}$4.14 $94.22_{1.37}$ $97.04_{0.89}$2.82
Llama-3.2 11B $79.88_{2.96}$ $68.64_{3.75}$ 11.24 $90.88_{1.56}$ $82.91_{2.11}$ 7.97
Llama-3.2 90B $79.29_{3.14}$ $71.01_{3.36}$ 8.28 $87.81_{2.00}$ $83.17_{2.30}$ 4.64
@@
Qwen2-VL 7B $90.53_{2.25}$ $\underline{\bm{96.45_{1.39}}}$5.92 $\bm{94.28_{1.54}}$ $\underline{\bm{98.82_{0.49}}}$4.54
Qwen2-VL 72B $94.08_{1.80}$ $95.27_{1.67}$1.18 $96.37_{1.23}$ $97.49_{0.98}$1.12
Qwen2.5-VL 7B $\underline{\bm{95.86_{1.53}}}$ $94.67_{1.76}$ 1.18 $\underline{\bm{98.52_{0.55}}}$ $96.51_{1.25}$ 2.01
Qwen2.5-VL 72B $93.49_{1.91}$ $90.53_{2.27}$ 2.96 $95.54_{1.41}$ $91.80_{2.05}$ 3.74
@@
CogVLM2-FT 19B $75.74_{0.72}$ $67.46_{0.64}$ 8.28 $\underline{\bm{90.6_{0.13}}}$ $84.26_{0.08}$ 6.34
@@
InternVL-V2 76B $22.34_{3.10}$ $23.94_{3.19}$1.60 $46.64_{2.46}$ $48.01_{2.51}$1.36
Llama-3.2 11B $10.65_{2.33}$ $7.69_{2.04}$ 2.96 $33.50_{2.28}$ $26.80_{2.03}$ 6.70
Llama-3.2 90B $13.02_{2.59}$ $15.38_{2.75}$2.37 $36.80_{2.36}$ $39.85_{2.52}$3.06
@@
Qwen2-VL 7B $\bm{75.74_{3.32}}$ $73.96_{3.55}$ 1.78 $85.91_{2.17}$ $85.83_{2.04}$ 0.08
Qwen2-VL 72B $71.60_{3.61}$ $72.19_{3.60}$0.59 $84.52_{2.18}$ $86.17_{1.89}$1.64
Qwen2.5-VL 7B $\underline{\bm{82.84_{2.91}}}$ $\underline{75.74_{3.30}}$ 7.10 $93.38_{1.21}$ $\underline{88.75_{1.73}}$ 4.64
Qwen2.5-VL 72B $81.07_{2.98}$ $74.56_{3.29}$ 6.51 $89.03_{2.01}$ $83.21_{2.52}$ 5.82
@@
CogVLM2-Chinese-FT 19B $66.49_{0.74}$ $67.55_{0.73}$1.06 $79.48_{0.17}$ $81.78_{0.09}$2.3
@@
Qwen2-VL 7B $67.55_{3.34}$ $73.40_{3.21}$5.85 $84.63_{1.80}$ $87.12_{1.76}$2.49
Qwen2-VL 72B $70.74_{3.46}$ $78.72_{3.00}$7.98 $85.14_{1.87}$ $90.40_{1.40}$5.26
Qwen2.5-VL 7B $73.40_{3.32}$ $53.19_{3.60}$ 20.21 $86.14_{1.81}$ $62.55_{3.20}$ 23.58
Qwen2.5-VL 72B $\underline{\bm{77.13_{3.02}}}$ $\underline{\bm{81.91_{2.87}}}$4.79 $\underline{\bm{88.81_{1.65}}}$ $\underline{\bm{90.67_{1.54}}}$1.87
@@
CogVLM2-Chinese-FT 19B $\underline{\bm{46.81_{0.32}}}$ $\underline{\bm{46.28_{0.49}}}$ 0.53 $\underline{\bm{66.85_{0.39}}}$ $\underline{\bm{69.79_{0.12}}}$2.95
@@
Table 9: Results of various open-source and closed-source vision language models on the VCR task using the first 500 test cases. FT = fine-tuned on 16,000 samples from the VCR-wiki training set. We label the best result of each setting and metric with bold fonts, and the best open-souce model with underline. Subscripts are standard deviations obtained from Bootstrap. For more latest models’ results, please visit our GitHub repository.
@@
CogVLM2-FT 19B $92.8_{0.06}$ $92.67_{0.13}$ 0.12 $97.51_{0.24}$ $97.45_{0.07}$ 0.06
@@
InternVL-V2 76B $81.24_{1.40}$ $90.06_{1.06}$8.82 $90.64_{0.77}$ $96.06_{0.46}$5.42
Llama-3.2 11B $79.25_{1.40}$ $66.46_{1.63}$ 12.80 $89.98_{0.77}$ $80.91_{1.06}$ 9.08
Llama-3.2 90B $80.87_{1.37}$ $71.55_{1.54}$ 9.32 $89.63_{0.85}$ $84.34_{0.98}$ 5.29
@@
Qwen2-VL 7B $90.06_{1.07}$ $94.53_{0.82}$4.47 $93.77_{0.76}$ $\underline{\bm{97.80_{0.34}}}$4.03
Qwen2-VL 72B $92.42_{0.92}$ $94.66_{0.78}$2.24 $94.77_{0.72}$ $97.48_{0.42}$2.71
Qwen2.5-VL 7B $\underline{\bm{94.91_{0.79}}}$ $\underline{\bm{93.79_{0.85}}}$ 1.12 $\underline{\bm{98.17_{0.29}}}$ $96.95_{0.47}$ 1.22
Qwen2.5-VL 72B $92.67_{0.91}$ $90.19_{1.08}$ 2.48 $95.55_{0.65}$ $92.52_{0.86}$ 3.02
@@
CogVLM2-FT 19B $75.9_{0.13}$ $65.22_{0.18}$ 10.68 $89.75_{0.14}$ $82.71_{0.27}$ 7.04
@@
InternVL-V2 76B $20.76_{1.29}$ $19.36_{1.24}$ 1.40 $45.41_{1.14}$ $43.65_{1.10}$ 1.76
Llama-3.2 11B $13.91_{1.25}$ $7.33_{0.94}$ 6.58 $35.78_{1.14}$ $26.14_{0.94}$ 9.64
Llama-3.2 90B $15.16_{1.28}$ $12.17_{1.13}$ 2.98 $37.57_{1.13}$ $35.14_{1.04}$ 2.43
@@
Qwen2-VL 7B $76.27_{1.49}$ $\underline{75.65_{1.45}}$ 0.62 $86.56_{1.00}$ $86.77_{0.93}$0.22
Qwen2-VL 72B $70.56_{1.68}$ $73.42_{1.53}$2.86 $82.94_{1.04}$ $86.69_{0.87}$3.74
Qwen2.5-VL 7B $80.12_{1.38}$ $74.04_{1.51}$ 6.09 $\underline{\bm{91.68_{0.66}}}$ $\underline{87.85_{0.81}}$ 3.83
Qwen2.5-VL 72B $\underline{\bm{80.50_{1.41}}}$ $70.06_{1.62}$ 10.43 $88.91_{0.88}$ $79.05_{1.25}$ 9.86
@@
CogVLM2-Chinese-FT 19B $63.97_{0.55}$ $62.67_{0.17}$ 1.3 $79.71_{0.41}$ $79.22_{0.47}$ 0.49
@@
Qwen2-VL 7B $61.08_{1.52}$ $68.16_{1.46}$7.09 $78.38_{0.96}$ $83.48_{0.84}$5.10
Qwen2-VL 72B $66.47_{1.50}$ $\underline{\bm{74.75_{1.40}}}$8.28 $81.70_{0.88}$ $\underline{\bm{87.35_{0.75}}}$5.66
Qwen2.5-VL 7B $72.16_{1.44}$ $47.21_{1.59}$ 24.95 $84.40_{0.83}$ $57.99_{1.34}$ 26.41
Qwen2.5-VL 72B $\underline{\bm{75.15_{1.41}}}$ $74.05_{1.37}$ 1.10 $\underline{\bm{86.54_{0.82}}}$ $84.29_{0.94}$ 2.25
@@
CogVLM2-Chinese-FT 19B $\underline{\bm{42.51_{0.32}}}$ $\underline{\bm{45.91_{0.23}}}$3.39 $\underline{\bm{65.79_{0.24}}}$ $\underline{\bm{69.46_{0.46}}}$3.68
@@
Appendix E More relationship between VCR and other benchmarks
@@
Appendix F Potential QA
@@
    Include VCR in VLM Pretraining: Just as OCR parsing tasks are often included in pretraining to improve OCR performance, researchers could consider incorporating VCR tasks during pretraining. We will codebase to facilitate this process, making it as straightforward as data augmentation.
@@
    Chain-of-Thought (CoT) Methods: Researchers could explore multi-modality pipelines based on CoT techniques to improve existing VLMs on VCR tasks (Chen et al., 2024a; Zhang et al., 2023). Although a model might not initially focus on the correct visual area (e.g., pixel-level hints in the TEI), CoT-based techniques could help refine its focus over successive rounds.
@@
Appendix G Real-world Case Study of Models Fine-tuned on VCR-wiki
[…]
Figure 7: Ground-truth:
BANK OF AMERICA TWO BRYANT PARK
[…]
Figure 8: Ground-truth:
METROPOLITAN
[…]
Figure 9: Ground-truth:
NOT IN SERVICE
[…]
Figure 10: Ground-truth:
NO CYCLING
[…]
Figure 11: Ground-truth:
SHIPPING FAX SERVICE
PASSPORT PHOTOS
COMPUTER RENTALS
[…]
Figure 12: Ground-truth:
Home of Peapack Private
[…]
Figure 13: Ground-truth:
DELICACY BREAKFAST LUNCH CATERING
[…]
Figure 14: Ground-truth:
INSPECT UPON RECEIPT…
DO NOT SIGN…
VISIBLE DAMAGE?
[…]
Figure 15: Ground-truth:
AMOUNT TIP TOTAL APPROVED
AMERICAN EXPRESS AID
THANK YOU / MERCI
CUSTOMER COPY
[…]
Figure 16: Ground-truth:
INSPECT UPON RECEIPT…
DO NOT SIGN…
VISIBLE DAMAGE?
[…]
Figure 17: Ground-truth:
AMOUNT TIP TOTAL APPROVED
AMERICAN EXPRESS AID
THANK YOU / MERCI
CUSTOMER COPY
[…]